# On robust overfitting: adversarial training induced distribution matters

### 28 Nov 2023 v1

@@ Appendix A Detailed Experimental setup
For PGD-AT, the settings on different datasets are summarized in Table 2. Data augmentation is performed on these datasets except for MNIST. during the training. For CIFAR-10 and CIFAR-100 we follow the data augmentation setting in [8]. For our reduced ImageNet, we adopt the same data augmentation scheme that is used on the restricted ImageNet in [49].
@@
For the IDEs on each datasets, the settings are outlined in Table 3. It is important to note that for each of the individual IDEs conducted on the same dataset, we maintain consistent training settings. This includes using the same model architecture with identical model size and the same level of regularization. This ensures a fair comparison of the IDE results obtained from the same dataset. Furthermore, the model is trained to achieve zero training error in all the IDEs, excluding the situation that the degeneration in model performance could be attributed to inadequate training procedures.
@@ Appendix B Proof of the theorem
We use the notations introduced in the main text. For shorter notations, let $z=(x,y)$ with $x$ following the distribution ${\cal D}_{\cal X}^{*}$ and $y=h^{*}(x)$. Let $u=(v,y)$ and $f(u):=f(v,y)$. We write $\mathcal{Q}_{x,y,\theta}$ as $\mathcal{Q}_{z}$, since our derivation does not explicitly depend on the choice of $\theta$.
@@
This is equivalent to saying that with probability $1-\tau$, we have
@@
For shorter notation, let $U:=(u_{1},\cdots,u_{m})$, $Z:=(z_{1},\cdots,z_{m})$, $\Gamma:=(\rho_{1},\cdots,\rho_{m})$, $F(Z,\Gamma):=\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho_{i}),y_{i})$. We have
@@
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\hat{Z}}\mathbb{E}_{\hat{\Gamma}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho_{i}),y_{i})-\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\hat{\rho}_{i}),\hat{y}_{i})\right|$ (25)
@@
$=$ $2\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho_{i}),y_{i})-\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho}_{i}),y_{i})\right|$ (31)
$=$ $\frac{2}{m}\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\mathbb{E}_{\Sigma}\left|\sum\limits_{i=1}^{m}\sigma_{i}\left(f(\mathcal{Q}_{z_{i}}(x_{i}+\rho_{i}),y_{i})-f(\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho}_{i}),y_{i})\right)\right|$ (32)
$\leq$ $\frac{2}{m}\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\sqrt{\sum\limits_{i=1}^{m}\left|f(\mathcal{Q}_{z_{i}}(x_{i}+\rho_{i}),y_{i})-f(\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho}_{i}),y_{i})\right|^{2}}$ (33)
$\leq$ $\frac{2}{m}\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\sqrt{\sum\limits_{i=1}^{m}\beta^{2}\|\mathcal{Q}_{z_{i}}(x_{i}+\rho_{i})-\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho}_{i})\|^{2}}$ (34)
$\leq$ $\frac{2\beta}{m}\mathbb{E}_{Z}\sqrt{\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\left[\sum\limits_{i=1}^{m}\|\mathcal{Q}_{z_{i}}(x_{i}+\rho_{i})-\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho}_{i})\|^{2}\right]}$ (35)
$=$ $\frac{2\beta}{m}\mathbb{E}_{Z}\sqrt{\sum\limits_{i=1}^{m}\mathbb{E}_{\rho}\mathbb{E}_{\bar{\rho}}\|\mathcal{Q}_{z_{i}}(x_{i}+\rho)-\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho})\|^{2}}$ (36)
$=$ $\frac{2\beta}{m}\mathbb{E}_{Z}\sqrt{\sum\limits_{i=1}^{m}\gamma(x_{i},y_{i})}$ (37)
$\leq$ $\frac{2\beta}{m}\sqrt{\mathbb{E}_{Z}\left[\sum\limits_{i=1}^{m}\gamma(x_{i},y_{i})\right]}$ (38)
$=$ $\frac{2\beta}{m}\sqrt{\sum\limits_{i=1}^{m}\mathbb{E}_{z_{i}}\gamma(x_{i},y_{i})}$ (39)
$=$ $\frac{2\beta}{\sqrt{m}}\sqrt{\mathbb{E}_{z}\gamma(x,y)}$ (40)
Again, we apply Jensen’s inequality to get (30). In (32), we introduce Rademacher variables $\Sigma:=(\sigma_{1},\cdots,\sigma_{m})$ (i.e., each random variable $\sigma_{i}$ takes values in $\{-1,+1\}$ independently with equal probability 0.5). The Rademacher variables introduces a random exchange of the corresponding difference term. Since $\Gamma$ and $\hat{\Gamma}$ are independently sampled from the same distribution, such a swap gives an equally likely configuration. Therefore, the equality (32) holds. The inequality (33) is given by Khintchine’s inequality. The inequality (34) makes use of the lipschitz condition of $f$. (35) is derived from Jensen’s inequality and due to that square root is a concave function. (37) is by the definition of the local dispersion of $\mathcal{Q}_{z}$. Again, we apply Jensen’s inequality to obtain (38). Equation (39) and (40) follow from the settings that each $z_{i}=(x_{i},y_{i})$ is i.i.d.
@@
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left|\mathbb{E}_{\bar{\Gamma}}\left[\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho}_{i}),y_{i})\right]-\mathbb{E}_{\tilde{\Gamma}}\left[\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\tilde{\rho}_{i}),\hat{y}_{i})\right]\right|$ (42)
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\bar{\rho}_{i}}\left[f(\mathcal{Q}_{z_{i}}(x_{i}+\bar{\rho}_{i}),y_{i})\right]-\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\tilde{\rho}_{i}}\left[f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\tilde{\rho}_{i}),\hat{y}_{i})\right]\right|$ (43)
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right]-\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right|$ (44)
$=$ $\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\mathbb{E}_{\Sigma}\left|\sum\limits_{i=1}^{m}\sigma_{i}\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|$ (45)
$\leq$ $\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\sqrt{\sum\limits_{i=1}^{m}\left|\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|^{2}}$ (46)
where equation (43) and (44) are due to each $\hat{\rho}_{i}$ and $\tilde{\rho}_{i}$ is i.i.d. Again, we introduce Rademacher variables at (45) and apply Khintchine’s inequality to get (46). For the term $\left|\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|^{2}$, we have
$\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})-\mathbb{E}_{\rho}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}$ (47)
$\leq$ $(\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right|+\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|)^{2}$ (48)
$\leq$ $2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right|^{2}+2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}$ (49)
@@
$\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\sqrt{\sum\limits_{i=1}^{m}\left|\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|^{2}}$
$\leq$ $\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\sqrt{\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right|^{2}+\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}}$ (50)
$\leq$ $\frac{1}{m}\sqrt{\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left[\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right|^{2}+\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}\right]}$ (51)
$=$ $\frac{1}{m}\sqrt{\sum\limits_{i=1}^{m}2\mathbb{E}_{z_{i}}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i}}(x_{i}+\rho),y_{i})\right|^{2}+\sum\limits_{i=1}^{m}2\mathbb{E}_{\hat{z}_{i}}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{\hat{z}_{i}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}}$ (52)
$=$ $\frac{2}{\sqrt{m}}\sqrt{\mathbb{E}_{z}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z}(x+\rho),y)\right|^{2}}$ (53)
$\leq$ $\frac{2}{\sqrt{m}}\sqrt{\sup\limits_{z\in{\rm supp}(\mathcal{D})}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z}(x+\rho),y)\right|^{2}}$ (54)
$=$ $\frac{2A}{\sqrt{m}}$ (55)
The final line is derived by the condition that $\sup\limits_{z\in{\rm supp}(\mathcal{D})}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z}(x+\rho),y)\right|=A$. This gives the final result
$\mathbb{E}_{U}g(u_{1}\cdots u_{m})\leq\frac{2\beta}{\sqrt{m}}\sqrt{\mathbb{E}_{z}\gamma(x,y)}+\frac{2A}{\sqrt{m}}$
Plugging back to (20), we derive the bound in Theorem 1. This completes the proof. $\Box$
Lastly we want to remark that the bound is not trivial, since we have
$A=$ $\sup\limits_{z\in{\rm supp}(\mathcal{D})}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z}(x+\rho),y)\right|$
$\leq$ $\sup\limits_{z\in{\rm supp}(\mathcal{D})}\mathbb{E}_{\rho}\left|f(\mathcal{Q}_{z}(x+\rho),y)\right|$
$\leq$ $\sup\limits_{z\in\mathcal{X}\times\mathcal{Y}}\sup\limits_{\|\rho\|_{\infty}\leq\epsilon}\left|f(\mathcal{Q}_{z}(x+\rho),y)\right|$
$\leq$ $\sup\limits_{v,y\in\mathcal{X}\times\mathcal{Y}}\left|f(v,y)\right|=B$
In fact, $A$ could be much smaller than $B$, meaning the bound is tight.

### 10 Feb 2024 v2

@@ Appendix A Detailed Experimental setup
For PGD-AT, the settings on different datasets are summarized in Table 2. Data augmentation is performed on these datasets except for MNIST during the training. For CIFAR-10 and CIFAR-100 we follow the data augmentation setting in [8]. For our reduced ImageNet, we adopt the same data augmentation scheme that is used on the restricted ImageNet in [50].
@@
For the IDEs on each datasets, the settings are outlined in Table 3. It is important to note that for each of the individual IDEs that is conducted on the same dataset, we maintain consistent training settings. This includes using the same model architecture with identical model size and the same level of regularization. This ensures a fair comparison of the IDE results obtained from the same dataset. Furthermore, the model is trained to achieve zero training error in all the IDEs, excluding the situation that the degeneration in model performance could be attributed to inadequate training procedures.
@@ Appendix B Proof of the theorem
We follow the notations introduced in the main text. Recall that $(x,y)$ denotes an instance drawn from the distribution ${\cal D}^{*}$ and $(v,y)$ denotes an instance drawn from the associate induced distribution $\tilde{\mathcal{D}}_{t}$. For shorter notations, we will denote $z:=(x,y)$, $u:=(v,y)$ and $f(u):=f(v,y)$ and write $\mathcal{Q}_{x,y,\theta_{t}}$ as $\mathcal{Q}_{z,\theta_{t}}$.
@@
where we use $U:=(u_{1},\cdots,u_{m})$. This is equivalent to saying that with probability $1-\tau$, we have
@@
For shorter notations, let $Z:=(z_{1},\cdots,z_{m})$, $\Gamma:=(\rho_{1},\cdots,\rho_{m})$ and $F(Z,\Gamma):=\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho_{i}),y_{i})$. We have
@@
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\hat{Z}}\mathbb{E}_{\hat{\Gamma}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho_{i}),y_{i})-\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{\hat{z}_{i},\theta_{t}}(\hat{x}_{i}+\hat{\rho}_{i}),\hat{y}_{i})\right|$ (23)
@@
$=$ $2\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho_{i}),y_{i})-\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho}_{i}),y_{i})\right|$ (29)
$=$ $\frac{2}{m}\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\mathbb{E}_{\Sigma}\left|\sum\limits_{i=1}^{m}\sigma_{i}\left(f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho_{i}),y_{i})-f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho}_{i}),y_{i})\right)\right|$ (30)
$\leq$ $\frac{2}{m}\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\sqrt{\sum\limits_{i=1}^{m}\left|f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho_{i}),y_{i})-f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho}_{i}),y_{i})\right|^{2}}$ (31)
$\leq$ $\frac{2}{m}\mathbb{E}_{Z}\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\sqrt{\sum\limits_{i=1}^{m}\beta^{2}\|\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho_{i})-\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho}_{i})\|^{2}}$ (32)
$\leq$ $\frac{2\beta}{m}\mathbb{E}_{Z}\sqrt{\mathbb{E}_{\Gamma}\mathbb{E}_{\bar{\Gamma}}\left[\sum\limits_{i=1}^{m}\|\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho_{i})-\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho}_{i})\|^{2}\right]}$ (33)
$=$ $\frac{2\beta}{m}\mathbb{E}_{Z}\sqrt{\sum\limits_{i=1}^{m}\mathbb{E}_{\rho}\mathbb{E}_{\bar{\rho}}\|\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho)-\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho})\|^{2}}$ (34)
$=$ $\frac{2\beta}{m}\mathbb{E}_{Z}\sqrt{\sum\limits_{i=1}^{m}\gamma_{t}(x_{i},y_{i})}$ (35)
$\leq$ $\frac{2\beta}{m}\sqrt{\mathbb{E}_{Z}\left[\sum\limits_{i=1}^{m}\gamma_{t}(x_{i},y_{i})\right]}$ (36)
$=$ $\frac{2\beta}{m}\sqrt{\sum\limits_{i=1}^{m}\mathbb{E}_{z_{i}}\gamma_{t}(x_{i},y_{i})}$ (37)
$=$ $\frac{2\beta}{\sqrt{m}}\sqrt{\mathbb{E}_{z}\gamma_{t}(x,y)}$ (38)
Again, we apply Jensen’s inequality to get (28). In (30), we introduce Rademacher variables $\Sigma:=(\sigma_{1},\cdots,\sigma_{m})$ (i.e., each random variable $\sigma_{i}$ takes values in $\{-1,+1\}$ independently with equal probability 0.5). The Rademacher variables introduces a random exchange of the corresponding difference term. Since $\Gamma$ and $\hat{\Gamma}$ are independently sampled from the same distribution, such a swap gives an equally likely configuration. Therefore, the equality (30) holds. The inequality (31) is given by Khintchine’s inequality. The inequality (32) makes use of the lipschitz condition of $f$. (33) is derived from Jensen’s inequality and due to that square root is a concave function. (35) is by the definition of the local dispersion of $\mathcal{Q}_{\theta_{t}}$. Again, we apply Jensen’s inequality to obtain (36). Equation (37) and (38) follow from the settings that each $z_{i}=(x_{i},y_{i})$ is i.i.d.
@@
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left|\mathbb{E}_{\bar{\Gamma}}\left[\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho}_{i}),y_{i})\right]-\mathbb{E}_{\tilde{\Gamma}}\left[\frac{1}{m}\sum\limits_{i=1}^{m}f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\tilde{\rho}_{i}),\hat{y}_{i})\right]\right|$ (40)
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\bar{\rho}_{i}}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\bar{\rho}_{i}),y_{i})\right]-\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\tilde{\rho}_{i}}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\tilde{\rho}_{i}),\hat{y}_{i})\right]\right|$ (41)
$=$ $\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left|\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right]-\frac{1}{m}\sum\limits_{i=1}^{m}\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right|$ (42)
$=$ $\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\mathbb{E}_{\Sigma}\left|\sum\limits_{i=1}^{m}\sigma_{i}\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|$ (43)
$\leq$ $\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\sqrt{\sum\limits_{i=1}^{m}\left|\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|^{2}}$ (44)
where equation (41) and (42) are due to each $\hat{\rho}_{i}$ and $\tilde{\rho}_{i}$ is i.i.d. Again, we introduce Rademacher variables at (43) and apply Khintchine’s inequality to get (44). For the term $\left|\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|^{2}$, we have
$\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})-\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}$ (45)
$\leq$ $(\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right|+\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|)^{2}$ (46)
$\leq$ $2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right|^{2}+2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}$ (47)
@@
$\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\sqrt{\sum\limits_{i=1}^{m}\left|\left(\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right]-\mathbb{E}_{\rho}\left[f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right]\right)\right|^{2}}$
$\leq$ $\frac{1}{m}\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\sqrt{\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right|^{2}+\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}}$ (48)
$\leq$ $\frac{1}{m}\sqrt{\mathbb{E}_{Z}\mathbb{E}_{\hat{Z}}\left[\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right|^{2}+\sum\limits_{i=1}^{m}2\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}\right]}$ (49)
$=$ $\frac{1}{m}\sqrt{\sum\limits_{i=1}^{m}2\mathbb{E}_{z_{i}}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(x_{i}+\rho),y_{i})\right|^{2}+\sum\limits_{i=1}^{m}2\mathbb{E}_{\hat{z}_{i}}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z_{i},\theta_{t}}(\hat{x}_{i}+\rho),\hat{y}_{i})\right|^{2}}$ (50)
$=$ $\frac{2}{\sqrt{m}}\sqrt{\mathbb{E}_{z}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z,\theta_{t}}(x+\rho),y)\right|^{2}}$ (51)
$\leq$ $\frac{2}{\sqrt{m}}\sqrt{\mathbb{E}_{z}(\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{z,\theta_{t}}(x+\rho),y)-f(x,y)\right|+\left|f(x,y)\right|)^{2}}$ (52)
$\leq$ $\frac{2}{\sqrt{m}}\sqrt{\mathbb{E}_{z}(\mathbb{E}_{\rho}\left|f(\mathcal{Q}_{z,\theta_{t}}(x+\rho),y)-f(x,y)\right|+B)^{2}}$ (53)
$\leq$ $\frac{2}{\sqrt{m}}\sqrt{\mathbb{E}_{z}(\mathbb{E}_{\rho}\beta\|\mathcal{Q}_{z,\theta_{t}}(x+\rho)-x\|_{2}+B)^{2}}$ (54)
$\leq$ $\frac{2(\beta\sqrt{d}\epsilon+B)}{\sqrt{m}}$ (55)
The inequalities (52)-(54) respectively make use of the triangle inequality, Jensen’s inequality, the boundedness and lipschitz condition of $f$. The final line is due to that with $\|\mathcal{Q}_{z,\theta_{t}}(x+\rho)-x\|_{\infty}\leq\epsilon$ we have $\|\mathcal{Q}_{z,\theta_{t}}(x+\rho)-x\|_{2}\leq\sqrt{d}\epsilon$. This gives the final result
$\mathbb{E}_{U}g(u_{1}\cdots u_{m})\leq\frac{2\beta}{\sqrt{m}}\sqrt{\mathbb{E}_{z}\gamma(x,y)}+\frac{2(\beta\sqrt{d}\epsilon+B)}{\sqrt{m}}$
Plugging back to (18), we derive the bound in the Theorem. $\Box$